OpenStudy (anonymous):

i need someone to EXPLAIN not give the answer. how would i simplify −2(n − 5)

OpenStudy (anonymous):

You have to distribute the -2 to n and -5

OpenStudy (anonymous):

in other words multiply -2 times n and -2 times -5

OpenStudy (mathstudent55):

You need to use the distributive property of multiplication over subtraction. The distributive property is the following: a(b - c) = a * b - a * c

OpenStudy (mathstudent55):

Now look at your problem and see how it follows the distributive property above: −2(n − 5) = -2 * n - (-2) * (5) = -2n - (-10) = -2n + 10